The standard mass of an unknown substance is 15.6g. Which of the following set has the least accurate value? show your solutions.


trial A B C

1 15.6 15.8 15.6

2 15.4 14.6 15.5

3 15.6 14.7 15.8


a. A b. B c. C d. D


1
Expert's answer
2021-09-24T13:25:14-0400

"(15.6+15.8+15.6)/3=15.67\ (g)"


"(15.4+14.6+15.5)/3=15.17\ (g)" "\sqrt{}"


"(15.6+14.7+15.8)/3=15.37\ (g)"



The second set has the least accurate.